# Average entropy of Gaussian mixtures

### v1

@@ 2.5 Pulling leading-order terms out of the integral
We now use Corollary 2.3 and Lemma 2.6 to further obtain an expression for the differential entropy that separates the leading contributions and the quantity $S$ which we define as follows:
@@
Let $h_{\sigma}=\ln\sigma\sqrt{2\pi e}$ be the differential entropy of a Gaussian distributed random variable with variance $\sigma$. The differential entropy $h({\mathbf{X}}|\hat{\mathbf{W}})$ is given by:
@@
2.6 Third change of variables; simplification
We would like to approximate $S$ given in equation (31). However, our current integration variables $\{\mathbf{u}_{k}\}_{k=1}^{q}$ do not give simple formulas for $G_{l}$, $T$, or $S$. By linearly mixing $\mathbf{u}_{q-1}$ and $\mathbf{u}_{q}$, we obtain a set of $q$ independent standard normal-distributed variables with the benefit that $G_{l}$ contains only one of the two newly introduced variables. We define the transformation matrix $W$ as:
@@ 2.6 Third change of variables; simplification
Let $\mathbf{r}=\sqrt{m_{2}}\mathbf{u}_{q}$, $\mathbf{s}=\sqrt{m_{1}}\mathbf{u}_{q-1}$, and $W$ as given in equation (34), then the variables $\tilde{\mathbf{r}}$ and $\tilde{\mathbf{s}}$ given by:
@@
have independent standard normal distributions, i.e. we can write:
@@
where $[\tilde{\cdot}]$ denotes the transformed expression, and $\mathbb{E}_{u,\tilde{s},\tilde{r}}[\tilde{\cdot}]$ is taken w.r.t the density
@@
then after applying the change of variables in Lemma 2.8, $T$ and $G_{l}$ can be written as:
@@
In this new form (38,39), the powers of $\mu$ are explicitly visible. The variables $\mathbf{w}_{l},\tilde{\mathbf{r}},\tilde{\mathbf{s}}$ do not generate any $\mu$-dependence when integrated. It is clear that $\mu G_{l}$ and $\mu T$ are of order $\mathcal{O}(\sqrt{\mu})$; hence we can formally proceed with a Taylor expansion of the $\mathrm{exp}$ function in (31).
@@ 3 Brute force expansion
We develop a series expansion in $\mu$ as follows. First we expand the exp function in (31), with $G_{l}$ and $T$ as defined in (38,39). This yields a complicated series, whose leading order behaviour is $1+{\cal O}(\sqrt{\mu})$. Then we substitute that series into the Taylor expansion of $\ln(1+\varepsilon)$ in order to evaluate $S$ as a series expansion. Finally we apply the expectation $\mathbb{E}_{u,\tilde{s},\tilde{r}}$, which is possible because we only get expectations of powers of $\mathbf{u},\tilde{\mathbf{r}},\tilde{\mathbf{s}}$ which are independent normal-distributed variables. Although all steps by themselves are straightforward enough, the result is a huge amount of bookkeeping. A number of things are worth noting:
  • What helps in this exercise is that odd powers of $\mathbf{w}_{l}$, $\tilde{\mathbf{r}}$ and $\tilde{\mathbf{s}}$ lead to vanishing integrals.
  • Furthermore, it is clear from the start that odd powers of $\sqrt{\mu}$ will disappear in the end. This can be seen from the fact that a factor $\frac{1}{\sqrt{\mu}}$ in $T$ and $G_{l}$ always occurs with an isolated $\tilde{\mathbf{r}}$; hence any occurrence of an odd power of $\sqrt{\mu}$ comes with an odd power of $\tilde{\mathbf{r}}$.
  • Due to the $n$-dimensional inner products that occur in $G_{l}$ and $T$, which consist of $n$ independent terms, each power of $\mu$ is associated with a factor $n$. Consequently, the power series becomes a series not in $\mu$ but actually in $n\mu$. For convergence, the product $n\mu$ needs to be sufficiently small. Fortunately we are allowed to work under the condition $n\leq q$, as explained in Section 2.2.
We have performed this brute-force exercise up to order $\mathcal{O}(\mu^{2})$.
@@
The differential entropy $h({\mathbf{X}}|\hat{\mathbf{W}})$ up to order $\mathcal{O}(\mu^{2})$ is given by:
@@
Theorem 3.1 is not directly comparable to results in the literature. However, we can make use of existing upper bounds in order to assess our expression. The Gaussian upper bound in equation (1) is known to be loose, while the bound in (4) is closer to the value of the differential entropy. For our case, the bound in equation (4) reads:
$h({\mathbf{X}}|\hat{\mathbf{W}})\leq h_{\sigma}+\ln q.$ (41)
It is easy to see that for $\mu\leq q^{-1}$ and $n=q$, our second order Taylor expansion in equation (40) satisfies this bound up to the error produced by the truncation of the series, which is of order $o(\mu^{2})$.
@@ 4 Determinant-based approach
The $\hat{P}$ is essentially a $2\times 2$ matrix. The $\hat{Q}_{\ell}$ consists of the projector $\hat{\alpha}_{\ell}\hat{\alpha}_{\ell}^{T}$ plus extra entries in row and column $q-1$. Let $Z$ be defined as:
$Z=1-q^{-1}\sum_{l=1}^{q}\exp\left\{-\frac{\mu}{2}\left(G_{l}+2\delta_{q,l}T\right)\right\}=1-q^{-1}\sum_{l=1}^{q-1}\exp\left\{-\frac{1}{2}\hat{\mathbf{a}}^{\mathrm{T}}\hat{Q}_{l}\hat{\mathbf{a}}\right\}-q^{-1}\exp\left\{-\frac{1}{2}\hat{\mathbf{a}}^{\mathrm{T}}\hat{P}\hat{\mathbf{a}}\right\},$ (49)
where equality follows from the definitions (45). We note that $Z={\cal O}(\sqrt{\mu})$. This allows us to do an expansion in $Z$, knowing that $Z^{k}$ does not produce powers of $\mu$ lower than $\mu^{k/2}$. Furthermore, the power $\sqrt{\mu}$ occurs only in the off-diagonal components of $\hat{P}$ and $\hat{Q}_{\ell}$, whereas the diagonal components are ${\cal O}(\mu)$. As in the brute-force approach, the integration over $\hat{\mathbf{a}}$ eliminates half-integer powers of $\mu$. Hence ${\mathbb{E}}_{a}Z^{k}={\cal O}(\mu^{\lceil k/2\rceil})$. In order to obtain the entropy estimation up to and including power $\mu^{t}$, we need all contributions up to and including ${\mathbb{E}}_{a}Z^{2t}$. Using $\ln(1-Z)=-\sum_{k=1}^{\infty}k^{-1}Z^{k}$, one can easily show that $S$ (43) can be written as:
@@
This provides a recipe for going to arbitrary order in $\mu$. We show the calculation of the $\mu^{1}$ contribution in $S$. For this we need only ${\cal Z}_{1}$ and ${\cal Z}_{2}$.
@@
For a positive definite matrix $\hat{V}$, we have $\int\exp\left\{-\dfrac{1}{2}\mathbf{\hat{u}}^{\mathrm{T}}\hat{V}\mathbf{\hat{u}}\right\}\mathrm{d}\mathbf{u}_{1}\cdots\mathrm{d}\mathbf{u}_{q}=\dfrac{(2\pi)^{nq/2}}{\det(\hat{V})^{n/2}}$.
@@
Using (33) it follows that $h({\mathbf{X}}|\hat{\mathbf{W}})=nh_{\sigma}+\frac{n}{2}\frac{q}{q-1}\mu+S$, with $S\approx{\cal Z}_{1}+{\cal Z}_{2}$, yielding $h({\mathbf{X}}|\hat{\mathbf{W}})=nh_{\sigma}+\frac{n}{2}(1-\frac{1}{q})\mu+{\cal O}(\mu^{2})$. This is consistent with the result of the brute-force approach, Theorem 3.1.
In higher orders of $Z$, determinants occur of the form $\det(\hat{I}+r\hat{P}+t\hat{Q}_{\ell}+t^{\prime}\hat{Q}_{\ell^{\prime}}+t^{\prime\prime}\hat{Q}_{\ell^{\prime\prime}}+\cdots)$ for integer $r,t,t^{\prime},t^{\prime\prime},\ldots$. These are computable but involve a lot of bookkeeping. Because of the form $[\det(\cdots)]^{-n/2}$ and the fact that each determinant starts with 1, followed by integer powers of $\mu$, the leading order coefficient of $\mu^{t}$ is proportional to $\binom{n/2}{t}\propto n^{t}$. Hence, as in the brute force approach, the expansion in powers of $\mu$ effectively becomes an expansion in powers of $n\mu$.
@@ 5 Discussion
Gaussian mixtures occur in the literature either as an approximation to a probability density function or as the actual density of a two-stage stochastic process. Shannon entropy plays an important role in signal processing, optimisation and analysis of systems. It is important to have (semi-)analytical methods for evaluating the Shannon entropy (differential entropy) of Gaussian mixtures, without lengthy numerical computations in high-dimensional spaces. We have developed an expansion method which has a number of advantages: (i) it yields an entropy estimate rather than merely bounds; (ii) the order of magnitude of the error is known; (iii) it avoids the process of [4] that splits wide Gaussians into narrow pieces, which leads to a large number of expansion terms and introduces inaccuracies. The inaccuracies in [4] grow with the number of wide Gaussians in the mixture. The approach in [4] is unable to exploit the existence of a small parameter $\mu$ when the mixture structure allows it. In our case, we make sure to leverage the trade-off between the support shared between the components and the variance of the components, which makes our method more efficient when the mixture is not composed of separated clusters. In the case of negligible shared support between the components, an approximation is not needed as the bound in equation (41) becomes exact. From the second order result (40) it seems that convergence is faster than what we expected. In leading order terms, every power of $\mu$ is accompanied by a power of $n$, which may endanger the radius of convergence. However, in the $\mu^{2}$ part in (40) we see that the additional power of $n$ comes as $n/q$, which is not dangerous since we are allowed to work with $n\leq q$. Setting $n=q$ and $q\gg 1$, (40) reduces to $h({\mathbf{X}}|\hat{\mathbf{W}})\approx\frac{n}{2}[\ln 2\pi e\sigma^{2}+\mu-\frac{1}{q}\mu^{2}+{\cal O}(\mu^{3})]$, which has fast convergence. In future work we will try to apply our method for fixed displacements $\hat{\mathbf{w}}$, and study the convergence at higher orders of $\mu$.
@@ Acknowledgements
Part of this work was supported by NWO grant CS.001 (Forwardt).
@@
All the integrals are from $\infty$- ∞ to $\infty$. Inside the $x$-integration we define new integration variables ${\mathbf{z}}_{j}$ by writing ${\mathbf{w}}_{j}={\mathbf{x}}-s{\mathbf{z}}_{j}$. Furthermore we write ${\mathbf{x}}=s{\mathbf{c}}$. Now we have:
@@
therefore for any $F$ that is permutation invariant, we can write:
@@
Suppose some eigenvectors of $\hat{C}^{(q)}$ are of the form $(1,\dots,1,x)^{\mathrm{T}}$ where $x$ is unknown. We now wish to solve the eigenvalue equation:
@@
for some eigenvalue $\lambda$. Carrying out the matrix multiplication, we get:
@@
From equation (66), get the two following values for $x$:
@@
where we write $m_{1}=\lambda_{1}+1$ and $m_{2}=\lambda_{2}+1$, and we define $\hat{x}_{1}$ and $\hat{x}_{2}$ by:
@@
Note that $\hat{x}_{1}$ and $\hat{x}_{2}$ are orthogonal, that is using equation (77) we have:
@@
We define the normalized eigenvectors $\hat{\beta}_{1}$ and $\hat{\beta}_{2}$ by:
@@
which in matrix notation are given by:
@@
One can check that $\hat{y}_{i}$ is an eigenvector of $\hat{C}^{(q)}$ with eigenvalue 0. The normalized eigenvectors $\hat{\alpha}_{i}$ are defined by:
@@
It can be checked that the set $\{\hat{\alpha}_{i}\}_{i=1}^{q-2}\cup\{\hat{\beta}_{1},\hat{\beta}_{2}\}$ are orthonormal. Therefore, we can form the orthogonal matrix $\hat{\Lambda}^{(q)}$ with its columns being eigenvectors of $\hat{C}^{(q)}$, that is:
@@
We perform the change of variables:
@@
Note that we can write:
@@
and note that $\alpha_{m,q}=0$, so we have:
@@
where note that $G_{l}=0$ if $l=q$, and so we have:
@@
Rearranging equation (D) further yields the desired expression for $T$. Writing $F$ in terms of the new variables, we get:
@@
We have $\mathrm{d}\mathbf{z}_{1}\cdots\mathrm{d}\mathbf{z}_{q}=\mathrm{d}\mathbf{u}_{1}\cdots\mathrm{d}\mathbf{u}_{q}$ since $\hat{\Lambda}^{(q)}$ is orthogonal, and we also have:
@@
It then follows directly from Lemma 2.6 that:
$\int\exp\left\{-\dfrac{1}{2}\mathbf{\hat{z}}^{\mathrm{T}}\hat{M}^{(q)}\mathbf{\hat{z}}\right\}F(\mathbf{z}_{1},\dots,\mathbf{z}_{q})\mathrm{d}\mathbf{z}_{1}\cdots\mathrm{d}\mathbf{z}_{q}=$ (106)
$-\dfrac{\mu}{2}\int\exp\left\{-\dfrac{1}{2}\mathbf{\hat{u}}^{\mathrm{T}}\hat{D}^{(q)}\mathbf{\hat{u}}\right\}\left(\dfrac{1}{||\hat{x}_{1}||^{2}}\mathbf{u}_{q-1}\cdot\mathbf{u}_{q-1}+\dfrac{1}{||\hat{x}_{2}||^{2}}\mathbf{u}_{q}\cdot\mathbf{u}_{q}+\dfrac{2}{||\hat{x}_{1}||||\hat{x}_{2}||}\mathbf{u}_{q-1}\cdot\mathbf{u}_{q}\right)\mathrm{d}\mathbf{u}_{1}\cdots\mathrm{d}\mathbf{u}_{q}$
@@
Note that we have:
@@
Rearranging equation (108), we have:
@@
where $g_{\tilde{\mu},\sigma^{2}}(x)$ is the Gaussian function with mean $\tilde{\mu}$ and variance $\sigma^{2}$. The first term in equation (106) yields:
@@
$=-\dfrac{\mu}{2||\hat{x}_{1}||^{2}}\sum_{j=1}^{n}\int u^{2}_{q-1,j}\prod_{i=1}^{n}\prod_{k=1}^{q}\dfrac{\sqrt{2\pi}}{\sqrt{d_{k,i}}}g_{0,d^{-1}_{k,i}}(u_{k,i})\mathrm{d}u_{k,i}=-\dfrac{\mu(\sqrt{2\pi})^{nq}}{2||\hat{x}_{1}||^{2}}\times$
$\sum_{j=1}^{n}\left(\int\prod_{(k,i)\neq(q-1,j)}\dfrac{1}{\sqrt{d_{k,i}}}g_{0,d^{-1}_{k,i}}(u_{k,i})\mathrm{d}u_{k,i}\right)\left(\dfrac{1}{\sqrt{d_{q-1,j}}}\int g_{0,d^{-1}_{q-1,j}}(u_{q-1,j})u^{2}_{q-1,j}\mathrm{d}u_{q-1,j}\right)$
$=-\dfrac{\mu(\sqrt{2\pi})^{nq}}{2||\hat{x}_{1}||^{2}}\sum_{j=1}^{n}\left(\prod_{(k,i)\neq(q-1,j)}\dfrac{1}{\sqrt{d_{k,i}}}\right)\left(\dfrac{d^{-1}_{q-1,j}}{\sqrt{d_{q-1,j}}}\right)$
$=-\dfrac{\mu(\sqrt{2\pi})^{nq}}{2||\hat{x}_{1}||^{2}}\sum_{j=1}^{n}\left(\prod_{i=1}^{n}\prod_{k=1}^{q}\dfrac{1}{\sqrt{d_{k,i}}}\right)d^{-1}_{q-1,j}=-\dfrac{\mu(\sqrt{2\pi})^{nq}}{2||\hat{x}_{1}||^{2}}\sum_{j=1}^{n}\dfrac{1}{\left(\sqrt{m_{1}m_{2}}\right)^{n}}\dfrac{1}{m_{1}}$
@@
In an analogous calculation, the second term in equation (106) yields:
@@
The third term in equation (106) is identically zero since the matrix $\hat{D}^{(q)}$ is diagonal (there is no covariance). From Corollary 2.3, equations (111), (112), (74), and (78), we have:

### v2

@@ 2.5 Pulling Leading-Order Terms Out of the Integral
We now use Corollary 2.3 and Lemma 2.6 to further obtain an expression for the differential entropy that separates the leading contributions and the quantity $S$, which we define as follows:
@@
Let $h_{\sigma}=\ln\sigma\sqrt{2\pi e}$ be the differential entropy of a Gaussian distributed random variable with variance $\sigma$. The differential entropy $h({\mathbf{X}}|\hat{\mathbf{W}})$ is given by
@@
2.6 Third Change of Variables: Simplification
We would like to approximate $S$ given in Equation (27). However, our current integration variables $\{\mathbf{u}_{k}\}_{k=1}^{q}$ do not give simple formulas for $G_{l}$, $T$, or $S$. By linearly mixing $\mathbf{u}_{q-1}$ and $\mathbf{u}_{q}$, we obtain a set of $q$-independent standard normal-distributed variables with the benefit that $G_{l}$ contains only one of the two newly introduced variables. We define the transformation matrix $W$ as
@@ 2.6 Third Change of Variables: Simplification
Let $\mathbf{r}=\sqrt{m_{2}}\mathbf{u}_{q}$, $\mathbf{s}=\sqrt{m_{1}}\mathbf{u}_{q-1}$, and $W$ be as given in Equation (30), then the variables $\tilde{\mathbf{r}}$ and $\tilde{\mathbf{s}}$ given by
@@
have independent standard normal distributions, i.e., we can write
@@
where $[\tilde{\cdot}]$ denotes the transformed expression, and $\mathbb{E}_{u,\tilde{s},\tilde{r}}[\tilde{\cdot}]$ is taken with regard to the density
@@
then after applying the change of variables in Lemma 2.8, $T$ and $G_{l}$ can be written as
@@
In this new form (34) and (35), the powers of $\mu$ are explicitly visible. The variables $\mathbf{w}_{l},\tilde{\mathbf{r}},\tilde{\mathbf{s}}$ do not generate any $\mu$-dependence when integrated. It is clear that $\mu G_{l}$ and $\mu T$ are of order $\mathcal{O}(\sqrt{\mu})$; hence, we can formally proceed with a Taylor expansion of the $\mathrm{exp}$ function in (27).
@@ 3 Brute Force Expansion
We develop a series expansion in $\mu$ as follows. First, we expand the exp function in (27), with $G_{l}$ and $T$ as defined in (34) and (35). This yields a complicated series, whose leading order behaviour is $1+{\cal O}(\sqrt{\mu})$. Then, we substitute that series into the Taylor expansion of $\ln(1+\varepsilon)$ in order to evaluate $S$ as a series expansion. Finally, we apply the expectation $\mathbb{E}_{u,\tilde{s},\tilde{r}}$, which is possible because we only obtain expectations of powers of $\mathbf{u},\tilde{\mathbf{r}},\tilde{\mathbf{s}}$, which are independent normal-distributed variables. Although all steps by themselves are straightforward enough, the result is a huge amount of bookkeeping. A number of things are worth noting:
  • What helps in this exercise is that odd powers of $\mathbf{w}_{l}$, $\tilde{\mathbf{r}}$, and $\tilde{\mathbf{s}}$ lead to vanishing integrals.
  • Furthermore, it is clear from the start that the odd powers of $\sqrt{\mu}$ will disappear in the end. This can be seen from the fact that a factor $\frac{1}{\sqrt{\mu}}$ in $T$ and $G_{l}$ always occurs with an isolated $\tilde{\mathbf{r}}$; hence, any occurrence of an odd power of $\sqrt{\mu}$ comes with an odd power of $\tilde{\mathbf{r}}$.
  • Due to the $n$-dimensional inner products that occur in $G_{l}$ and $T$, which consist of $n$ independent terms, each power of $\mu$ is associated with a factor $n$. Consequently, the power series becomes a series not in $\mu$ but actually in $n\mu$. For convergence, the product $n\mu$ needs to be sufficiently small. Fortunately, we are allowed to work under the condition $n\leq q$, as explained in Section 2.2.
We performed this brute-force exercise up to order $\mathcal{O}(\mu^{2})$.
@@
The differential entropy $h({\mathbf{X}}|\hat{\mathbf{W}})$ up to order $\mathcal{O}(\mu^{2})$ is given by
@@
Theorem 3.1 is not directly comparable with results in the literature. However, we can make use of existing upper bounds in order to assess our expression. The Gaussian upper bound in Equation (2) is known to be loose, while the bound in (5) is closer to the value of the differential entropy. For our case, the bound in Equation (5) reads
$h({\mathbf{X}}|\hat{\mathbf{W}})\leq nh_{\sigma}+\ln q.$ (37)
It is easy to see that for $\mu\leq q^{-1}$ and $n=q$, our second-order Taylor expansion in Equation (36) satisfies this bound up to the error produced by the truncation of the series, which is of order $o(\mu^{2})$.
@@ 4 Determinant-Based Approach
The $\hat{P}$ is essentially a $2\times 2$ matrix. The $\hat{Q}_{\ell}$ consists of the projector $\hat{\alpha}_{\ell}\hat{\alpha}_{\ell}^{T}$ plus extra entries in row and column $q-1$. Let $Z$ be defined as
$Z$ $=1-q^{-1}\sum_{l=1}^{q}\exp\left\{-\frac{\mu}{2}\left(G_{l}+2\delta_{q,l}T\right)\right\}$ (45)
$=1-q^{-1}\sum_{l=1}^{q-1}\exp\left\{-\frac{1}{2}\hat{\mathbf{a}}^{\mathrm{T}}\hat{Q}_{l}\hat{\mathbf{a}}\right\}-q^{-1}\exp\left\{-\frac{1}{2}\hat{\mathbf{a}}^{\mathrm{T}}\hat{P}\hat{\mathbf{a}}\right\},$
where equality follows from the definitions (41). We note that $Z={\cal O}(\sqrt{\mu})$. This allows us to perform an expansion in $Z$, knowing that $Z^{k}$ does not produce powers of $\mu$ lower than $\mu^{k/2}$. Furthermore, the power $\sqrt{\mu}$ occurs only in the off-diagonal components of $\hat{P}$ and $\hat{Q}_{\ell}$, whereas the diagonal components are ${\cal O}(\mu)$. As in the brute-force approach, the integration over $\hat{\mathbf{a}}$ eliminates half-integer powers of $\mu$. Hence, ${\mathbb{E}}_{a}Z^{k}={\cal O}(\mu^{\lceil k/2\rceil})$. In order to obtain the entropy estimation up to and including power $\mu^{t}$, we need all contributions up to and including ${\mathbb{E}}_{a}Z^{2t}$. Using $\ln(1-Z)=-\sum_{k=1}^{\infty}k^{-1}Z^{k}$, one can easily show that $S$ (39) can be written as
@@
This provides a recipe for going to arbitrary order in $\mu$. We show the calculation of the $\mu^{1}$ contribution in $S$. For this, we need only ${\cal Z}_{1}$ and ${\cal Z}_{2}$.
@@
For a positive definite matrix $\hat{V}$, we have
$\int\exp\left\{-\dfrac{1}{2}\mathbf{\hat{u}}^{\mathrm{T}}\hat{V}\mathbf{\hat{u}}\right\}\mathrm{d}\mathbf{u}_{1}\cdots\mathrm{d}\mathbf{u}_{q}=\dfrac{(2\pi)^{nq/2}}{\det(\hat{V})^{n/2}}$.
@@
Using (29) it follows that $h({\mathbf{X}}|\hat{\mathbf{W}})=nh_{\sigma}+\frac{n}{2}\frac{q}{q-1}\mu+S$, with $S\approx{\cal Z}_{1}+{\cal Z}_{2}$, yielding $h({\mathbf{X}}|\hat{\mathbf{W}})=nh_{\sigma}+\frac{n}{2}(1-\frac{1}{q})\mu+{\cal O}(\mu^{2})$. This is consistent with the result of the brute-force approach, as in Theorem 3.1. In higher orders of $Z$, determinants occur of the form $\det(\hat{I}+r\hat{P}+t\hat{Q}_{\ell}+t^{\prime}\hat{Q}_{\ell^{\prime}}+t^{\prime\prime}\hat{Q}_{\ell^{\prime\prime}}+\cdots)$ for integer $r,t,t^{\prime},t^{\prime\prime},\ldots$. These are computable but involve a lot of bookkeeping. Because of the form $[\det(\cdots)]^{-n/2}$ and the fact that each determinant starts with 1, followed by integer powers of $\mu$, the leading order coefficient of $\mu^{t}$ is proportional to $\binom{n/2}{t}\propto n^{t}$. Hence, as in the brute force approach, the expansion in powers of $\mu$ effectively becomes an expansion in powers of $n\mu$.
@@ 5 Discussion
Gaussian mixtures occur in the literature either as an approximation to a probability density function or as the actual density of a two-stage stochastic process. Shannon entropy plays an important role in the signal processing, optimization, and analysis of systems. It is important to have (semi-)analytical methods for evaluating the Shannon entropy (differential entropy) of Gaussian mixtures, without lengthy numerical computations in high-dimensional spaces. We have developed an expansion method that has a number of advantages: (i) it yields an entropy estimate rather than merely bounds; (ii) the order of magnitude of the error is known; (iii) it avoids the process of [7] that splits wide Gaussians into narrow pieces, which leads to a large number of expansion terms and introduces inaccuracies. The inaccuracies in [7] grow with the number of wide Gaussians in the mixture. The approach in [7] is unable to exploit the existence of a small parameter $\mu$ when the mixture structure allows it. In our case, we make sure to leverage the trade-off between the support shared between the components and the variance of the components, which makes our method more efficient when the mixture is not composed of separated clusters. In the case of negligible shared support between the components, an approximation is not needed as the bound in Equation (37) becomes exact. From the second-order result (36), it seems that convergence is faster than what we expected. In leading order terms, every power of $\mu$ is accompanied by a power of $n$, which may endanger the radius of convergence. However, in the $\mu^{2}$ part in (36), we see that the additional power of $n$ comes as $n/q$, which is not dangerous since we are allowed to work with $n\leq q$. Setting $n=q$ and $q\gg 1$, (36) reduces to $h({\mathbf{X}}|\hat{\mathbf{W}})\approx\frac{n}{2}[\ln 2\pi e\sigma^{2}+\mu-\frac{1}{q}\mu^{2}+{\cal O}(\mu^{3})]$, which has fast convergence. In future work, we will try to apply our method for fixed displacements $\hat{\mathbf{w}}$ and study the convergence at higher orders of $\mu$.
@@ Acknowledgements
This work was supported by NWO grant CS.001 (Forwardt) and by the Dutch Groeifonds project Quantum Delta NL KAT-2.
@@
All the integrals are from $\infty$- ∞ to $\infty$. Inside the $x$-integration, we define new integration variables ${\mathbf{z}}_{j}$ by writing ${\mathbf{w}}_{j}={\mathbf{x}}-s{\mathbf{z}}_{j}$. Furthermore, we write ${\mathbf{x}}=s{\mathbf{c}}$. Now, we have
@@
therefore, for any $F$ that is permutation invariant, we can write
@@
Suppose some eigenvectors of $\hat{C}^{(q)}$ are of the form $(1,\dots,1,x)^{\mathrm{T}}$, where $x$ is unknown. We now wish to solve the eigenvalue equation:
@@
for some eigenvalue $\lambda$. Carrying out the matrix multiplication, we obtain
@@
From Equation (62), obtain the two following values for $x$:
@@
where we write $m_{1}=\lambda_{1}+1$ and $m_{2}=\lambda_{2}+1$, and we define $\hat{x}_{1}$ and $\hat{x}_{2}$ by
@@
Note that $\hat{x}_{1}$ and $\hat{x}_{2}$ are orthogonal, that is, using Equation (73), we have
@@
We define the normalized eigenvectors $\hat{\beta}_{1}$ and $\hat{\beta}_{2}$ by
@@
which in matrix notation are given by
@@
One can check that $\hat{y}_{i}$ is an eigenvector of $\hat{C}^{(q)}$ with eigenvalue 0. The normalized eigenvectors $\hat{\alpha}_{i}$ are defined by
@@
It can be checked that the set $\{\hat{\alpha}_{i}\}_{i=1}^{q-2}\cup\{\hat{\beta}_{1},\hat{\beta}_{2}\}$ are orthonormal. Therefore, we can form the orthogonal matrix $\hat{\Lambda}^{(q)}$ with its columns being eigenvectors of $\hat{C}^{(q)}$, that is,
@@
We perform the following change of variables:
@@
Note that we can write
@@
and note that $\alpha_{m,q}=0$, so we have
@@
where note that $G_{l}=0$ if $l=q$, and so we have
@@
Rearranging Equation (D) further yields the desired expression for $T$. Writing $F$ in terms of the new variables, we obtain
@@
We have $\mathrm{d}\mathbf{z}_{1}\cdots\mathrm{d}\mathbf{z}_{q}=\mathrm{d}\mathbf{u}_{1}\cdots\mathrm{d}\mathbf{u}_{q}$ since $\hat{\Lambda}^{(q)}$ is orthogonal, and we also have
@@
It then follows directly from Lemma 2.6 that
$\int\exp\left\{-\dfrac{1}{2}\mathbf{\hat{z}}^{\mathrm{T}}\hat{M}^{(q)}\mathbf{\hat{z}}\right\}F(\mathbf{z}_{1},\dots,\mathbf{z}_{q})\mathrm{d}\mathbf{z}_{1}\cdots\mathrm{d}\mathbf{z}_{q}=-\dfrac{\mu}{2}\int\exp\left\{-\dfrac{1}{2}\mathbf{\hat{u}}^{\mathrm{T}}\hat{D}^{(q)}\mathbf{\hat{u}}\right\}\cdot$ (97)
$\left(\dfrac{1}{||\hat{x}_{1}||^{2}}\mathbf{u}_{q-1}\cdot\mathbf{u}_{q-1}+\dfrac{1}{||\hat{x}_{2}||^{2}}\mathbf{u}_{q}\cdot\mathbf{u}_{q}+\dfrac{2}{||\hat{x}_{1}||||\hat{x}_{2}||}\mathbf{u}_{q-1}\cdot\mathbf{u}_{q}\right)\mathrm{d}\mathbf{u}_{1}\cdots\mathrm{d}\mathbf{u}_{q}$
@@
Note that we have
@@
Rearranging Equation (99), we have
@@
where $g_{\tilde{\mu},\sigma^{2}}(x)$ is the Gaussian function with mean $\tilde{\mu}$ and variance $\sigma^{2}$. The first term in Equation (97) yields
@@
$=-\dfrac{\mu}{2||\hat{x}_{1}||^{2}}\sum_{j=1}^{n}\int u^{2}_{q-1,j}\prod_{i=1}^{n}\prod_{k=1}^{q}\dfrac{\sqrt{2\pi}}{\sqrt{d_{k}}}g_{0,d^{-1}_{k}}(u_{k,i})\mathrm{d}u_{k,i}=-\dfrac{\mu(\sqrt{2\pi})^{nq}}{2||\hat{x}_{1}||^{2}}\times$
$\sum_{j=1}^{n}\left(\int\prod_{(k,i)\neq(q-1,j)}\dfrac{1}{\sqrt{d_{k}}}g_{0,d^{-1}_{k}}(u_{k,i})\mathrm{d}u_{k,i}\right)\left(\dfrac{1}{\sqrt{d_{q-1}}}\int g_{0,d^{-1}_{q-1}}(u_{q-1,j})u^{2}_{q-1,j}\mathrm{d}u_{q-1,j}\right)$
$=-\dfrac{\mu(\sqrt{2\pi})^{nq}}{2||\hat{x}_{1}||^{2}}\sum_{j=1}^{n}\left(\prod_{(k,i)\neq(q-1,j)}\dfrac{1}{\sqrt{d_{k}}}\right)\left(\dfrac{d^{-1}_{q-1}}{\sqrt{d_{q-1}}}\right)$
$=-\dfrac{\mu(\sqrt{2\pi})^{nq}}{2||\hat{x}_{1}||^{2}}\sum_{j=1}^{n}\left(\prod_{i=1}^{n}\prod_{k=1}^{q}\dfrac{1}{\sqrt{d_{k}}}\right)d^{-1}_{q-1}=-\dfrac{\mu(\sqrt{2\pi})^{nq}}{2||\hat{x}_{1}||^{2}}\sum_{j=1}^{n}\dfrac{1}{\left(\sqrt{m_{1}m_{2}}\right)^{n}}\dfrac{1}{m_{1}}$
@@
In an analogous calculation, the second term in Equation (97) yields
@@
The third term in Equation (97) is identically zero since the matrix $\hat{D}^{(q)}$ is diagonal (there is no covariance). From Corollary 2.3, Equations (70), (74), (102) and (103), we have